Question 11 4 pts Assume that labor demand in a particular market continues to be given by H-800-20W, where H is measured in hours and W is in dollars. Suppose the government introduces subsidies to some lower-income households, leading to a labor supply curve of H 5W - 50 (instead of the previous H 5W). Given the new supply curve, the equilibrium number of hours worked is H hours. The equilibrium wage is W-$ (enter only numbers in the blanks, and round to the nearest integer if necessary).
Answer #1


Labor market is in equilibrium where the labor demand equals the labor supply.

Labor demand is as follows -

H = 800 - 20W

Labor supply is as follows -

H = 5W - 50

Equating the labor demand and the labor supply -

800 - 20W = 5W - 50

-20W - 5W = -50 - 800

-25W = -850

25W = 850

W = 850/25 = 34

H = 800 - 20W = 800 - (20*34) = 800 - 680 = 120 hours

So,

The equilibrium number of hours worked is H = 120 hours

The equilibrium wage is W = $34